Find the mean and the standard deviation of the heights of male engineering students in which the heights 66 inches and 75 inches correspond to standard score of -0.6 and 1.1, respectively.
1
Expert's answer
2020-11-05T11:49:44-0500

"\frac{66-\mu}{\sigma}=-0.6" and

"\frac{75-\mu}{\sigma}=1.1"

From the first equation "\frac{-(66-\mu)}{0.6}=\sigma"

"\frac{75-\mu}{\frac{-(66-\mu)}{0.6}}=1.1"

"{75-\mu}=1.1(\frac{-(66-\mu)}{0.6})"

"45-0.6\mu=1.1\mu-72.6"

"45+72.6=1.7\mu"

"\mu=\frac{117.6}{1.7}=" 69.176

"\frac{-(66-69.176)}{0.6}=\sigma"

"\sigma=5.294"